Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability

Expert replies
by mariah » Sat Nov 13, 2010 5:20 am
we have six color TV and two black-white TV. If two sets are to be selected at random, what is the probability that one black-white set is selected?



sorry don't have answers and oa
Join the discussion
Source: — Problem Solving |

by joshi.v123 » Sat Nov 13, 2010 5:29 am
is it 12/28?
Thanks,
VJ
Join the discussion

by mariah » Sat Nov 13, 2010 6:08 am
13/28 would be the answer if question asked what would b the probability if at least one black - white stem is selected...

but i think that this question asks probability of getting one Black - white and one color set (or vice versa) ...
Join the discussion

by beat_gmat_09 » Sat Nov 13, 2010 6:50 am
mariah wrote:we have six color TV and two black-white TV. If two sets are to be selected at random, what is the probability that one black-white set is selected?



sorry don't have answers and oa
two ways of getting -
BW C or C BW.
prob BW C = 2/8*6/7
prob C BW = 6/8*2/7
req prob - 2/8*6/7 + 6/8*2/7 = 3/7
Hope is the dream of a man awake
Join the discussion

by ov25 » Sat Nov 13, 2010 7:25 am
1-30/56 = 26/56 = 13/28

thanks
Join the discussion

by Deepthi Subbu » Tue Nov 16, 2010 9:12 pm
Am unable to understand , can somebody explain it further?
Join the discussion

by Geva@EconomistGMAT » Wed Nov 17, 2010 2:56 am
Beat_gmat_09 is correct.

Break the problem down to 1st choice, 2nd choice, find the probability of success in each, then multiply.

It seems as if the answer is P(BW on 1st) * P(color on 2nd)) = 2/8 * 6/7 = 1/2 * 3/7 = 3/14

However, that is only one of two possible scenarios of getting "one B/W", the other scenario being P(color on 1st) * P(B/W on 2nd) = 6/8 * 2/7 = 3/4 * 2/7 = 3/2 * 1/7 = 3/14. Not surprising the probability for both scenarios comes down to the same: there really is no reason why the probability of choosing a B/W first should be any different from the probability of choosing a B/W 2nd.

Since the final probability is made of P(choosing B/W on 1st) OR P(choosing B/W on 2nd), the final answer is 3/14+3/14 = 6/14 = 3/7.
Geva
Senior Instructor
Master GMAT
1-888-780-GMAT
https://www.mastergmat.com
Join the discussion